OpenStudy (anonymous):

suppose that two objects exert a gravitational force of 22 N on one another. what would that force be if the mass of one of the objects were doubled?

OpenStudy (mos1635):

F1=22 F1=G*M1*m2/r^2 F2=G*M1*2m2/r^2 do you see it?